wildhorse for life produces yoga mats. Mat a sells for $69 and has a contribution margin ratio of 40%. mat b sells for $122 and has a contribution margin ratio of 60% this year the company sold a total of 97600 mats of which 36600 where its of mat A. at the breakeven point the company needs to sell 31110 units of Mat A. how many units of Mat B were sold and what are the company's fixed costs?
Answer #1

QUESTION:-HOW MANY UNITS OF MAT B SOLD ?

ANSWER:- TOTAL NUMBER OF MATS SOLD 97600 MATS

NUMBER OF MAT A SOLD 36600 MATS

UNITS OF MAT B SOLD =97600-36600=61000

QUESTION:-COMPANY FIXED COST ?

ANSWERS:-FIXED COST=BREAK EVEN SALES*CONTRIBUTION MARGIN PERCENTAGE

=(31110*69)*40%

=$858636

AS STATED IN QUESTION COMPANYS' BREAKEVEN POINT IS WHEN 311110 UNITS OF MAT A IS SOLD SO PUTTING ABOVE FORMULA WE GOT THE COMPANY FIXED COST THAT IS $858636